# The Illusion of Readiness: Stress Testing Large Frontier Models on Multimodal Medical Benchmarks

### v2

@@ 2 Hidden Fragilities: Stress-Testing Models on Multimodal Medical benchmarks
Stress Test 6: Reasoning Signal
Objective: Characterize the reasoning behavior of multimodal models on multimodal medical benchmarks.
Setup: We analyze models from two perspectives. First, we apply Chain-of-Thought (CoT) [11] prompting to around 100 samples from NEJM and VQA-RAD [12] to test whether explicitly requesting step-by-step reasoning enhances performance. Second, we manually audit model-generated explanations across tasks, examining their factuality, visual grounding, and alignment with final answers. We also vary reasoning strength (from low to high) on OmniMedVQA [13] using OpenAI-o3 to assess its impact on reasoning quality.
Results: Prompting for CoT yields mixed or negative gains. On VQA-RAD and NEJM, CoT consistently reduces accuracy across models, except o4-mini (+2–5 pp), likely because its smaller capacity benefits from the explicit reasoning scaffolding that CoT provides. Similarly, increasing reasoning strength on OmniMedVQA offers only minor improvements without consistent benefit.
Manual audits reveal key reasoning failure patterns: Correct answers, incorrect logic: Models often produce plausible-sounding but factually wrong justifications—e.g., describing nonexistent features in the image or hallucinating visual findings. Reinforced visual misunderstanding: When initial visual interpretation is flawed, reasoning often amplifies the error rather than correcting it. Interpretable but uninformative steps: Some models generate structured reasoning steps (e.g., describing regions of the image), but these do not consistently guide the model to the correct answer. See Fig 4 for illustrative examples and Appendix C for additional, distinct cases.
Interpretation: Model-generated reasoning in medical benchmarks is often disconnected from true multimodal understanding. Correct predictions may result from pattern matching rather than reasoning, while rationales may confidently justify hallucinated or incorrect content. Even when visual reasoning appears structured, it frequently lacks functional linkage to the final answer. These findings underscore the unreliability of current reasoning signals and motivate a deeper examination of benchmark reasoning demands, which we present in Section 3.
Robustness Score
For each stress test (T1–T5), we computed a normalized robustness score in the range [0,1], with higher values indicating greater stability under perturbation. The mean robustness score was obtained by averaging across the five tests, providing a single measure of how reliably each model withstands missing modalities, format changes, distractor manipulations, and visual substitutions (Fig. 1). Full computational details are provided in Appendix A.
3 Benchmarking the Benchmarks: What Are We Really Measuring?
@@
Despite similar leaderboard scores, Section 2 revealed substantial variation in how models respond to medical stressors—ranging from hallucinated justifications to modality overreliance and format-based shortcuts. These divergent behaviors raise a fundamental question:
@@
If models succeed on benchmarks but fail under stress, what do these benchmarks actually measure—and how can we improve them?
@@
I. Motivation: Stress Test Failures Reveal Benchmark Gaps
The stress tests in Section 2 revealed that high benchmark scores do not guarantee robust model behavior. Models that performed well under standard conditions often failed when subjected to small perturbations—such as removing the image, reordering answer options, replacing distractors, or introducing misleading visual inputs.
Critically, these failure modes varied by benchmark. On NEJM, removing visual input led to steep performance drops, suggesting strong dependence on image understanding. In contrast, scores on JAMA remained stable—indicating that many items could be answered from text alone. Similarly, while chain-of-thought prompting provides explicit reasoning scaffolding that benefits distilled models such as o4-mini, its impact on NEJM was substantially greater than on VQA-RAD, underscoring the heterogeneous reasoning demands across benchmarks. In Test 5, even subtle visual substitutions caused top models to confidently choose incorrect answers—highlighting shortcut reliance on visual–answer associations in some datasets but not others.
@@ I. Motivation: Stress Test Failures Reveal Benchmark Gaps
II. Clinician-Informed Benchmark Profiling
To better understand what each benchmark actually tests, we developed a structured, clinician-guided rubric spanning ten medically meaningful dimensions. These axes capture key demands surfaced by our stress tests—ranging from reasoning complexity and medical context to uncertainty handling, visual detail, and multi-view alignment. The full rubric is shown in Fig 6.
Each benchmark was independently annotated by three board-certified clinicians per axis, using a 3-point ordinal scale. We report the median score for each benchmark-axis pair, and measure inter-annotator agreement using Fleiss’ $\kappa$ [19]. Agreement was moderate to strong across axes (range: 0.67–0.90). Highest agreement was observed on binary-seeming attributes such as Text-only Solvable ($\kappa$ = 0.90) and Clinical Context ($\kappa$ = 0.86); lower agreement appeared on more subjective dimensions such as Visual Detail Required ($\kappa$ = 0.67).
III. Visualizing the Benchmark Landscape
@@
  • VQA-RAD, PMC-VQA [17] and MIMIC-CXR [14] are visually dependent but low in inference complexity.
@@
This landscape helps to explain the benchmark-specific failures observed in Section 2. For instance, models trained on VQA-RAD may excel at image localization but struggle on diagnosis-oriented tasks like NEJM, which require cross-modal reasoning and medical inference.
Figure 5: Benchmarks differ in reasoning and visual complexity. Clinician-labeled profiles show that widely used multimodal benchmarks vary substantially in the reasoning and visual demands they place on models. While datasets like NEJM require high visual and inferential complexity, others like OmniMedVQA cluster low on both axes. These structural differences help explain the benchmark-specific fragilities observed in Section 2, and highlight the risk of treating all benchmarks as interchangeable measures of readiness.
IV. Implications for Model Evaluation and Design
These findings underscore the need to treat benchmarks not as monolithic performance targets, but as diagnostic tools that reveal distinct capabilities. We outline three implications:
@@ IV. Implications for Model Evaluation and Design
4 Rethinking Progress: Towards Real-world Readiness
@@
Across previous sections, we have shown that current benchmarks often overstate model readiness. Despite impressive leaderboard scores, many leading models exhibit inconsistent behavior, reliance on superficial patterns, and fragile reasoning under mild input perturbations. These findings challenge the prevailing assumption that benchmark success signals medical viability.
I. The False Signal of Benchmark Scores
@@
Today’s benchmarks offer a convenient but incomplete lens into model capability. Their design often emphasizes answer correctness while ignoring whether the answer is reached through medically valid reasoning, multimodal understanding, or robust generalization. As demonstrated in our stress tests, models can perform well by leveraging shallow cues—such as positional biases in answer options or memorized response formats—without engaging meaningfully with the underlying medical content.
Such behaviors are not easily detected by conventional metrics. For example, reasoning models may improve benchmark scores while still producing hallucinated or implausible rationales. Reinforcement learning approaches may optimize for token-level reward signals rather than faithful reasoning. The result is a misleading sense of progress: performance improves numerically, but not behaviorally.
II. Redefining Readiness in the Medical Context
Medical readiness is a multidimensional construct. In real-world settings, models must tolerate missing or noisy data, justify their decisions in a manner clinicians can understand, and reason across time, modality, and context. Performance must be not only accurate but also reliable, interpretable, and safe under uncertainty.
@@ II. Redefining Readiness in the Medical Context
Unlike domains such as mathematics or programming, where ground truth is unambiguous and logic is formal, healthcare requires flexible, contextual reasoning. Attempts to port reasoning strategies from other domains—such as multi-agent planning or chain-of-thought prompting—often fall short in medicine, where ambiguity, incomplete information, and medical nuance are the norm. Without alignment to this reality, advances in model capability risk being confined to artificial test settings.
III. Stress Testing as a Foundation for Trustworthy Evaluation
@@
Static benchmark scores provide little insight into how models behave under real-world uncertainty. To address this limitation, we propose structured stress testing as a core component of evaluation. As shown in Section 2, our tests T1–T5 revealed distinct vulnerabilities—models often succeed for the wrong reasons, retaining high accuracy even when visual input is removed (T1–T2, see Example  7-9); they show brittle behavior under minor perturbations such as distractor reordering, weakening, or visual substitution (T3–T5, See Example  7-12); and they frequently produce fabricated reasoning chains, generating confident but incorrect explanations that mask underlying shortcut behaviors (T6, See Example  13-15). These fragilities were hidden in leaderboard performance alone.
We generalize this approach into a framework of modular tests, each targeting a specific vulnerability such as neglect of visual input, dependence on spurious answer patterns, or overconfident reasoning when evidence is absent. By isolating these behaviors, stress tests enable systematic comparison across models and tasks, offering a principled means of diagnosing brittleness that static benchmarks obscure. In doing so, they establish a foundation for evaluation protocols that move beyond leaderboards and toward measures of robustness and trustworthiness.
IV. Toward Community Standards for Evaluation Reform
We propose several actionable recommendations for the field:
First, benchmarks should be accompanied by metadata that characterizes their reasoning and visual complexity, using structured rubrics like those presented in Section 3. This allows model performance to be interpreted in medically meaningful dimensions rather than aggregated scores alone.
Second, evaluation protocols should routinely incorporate adversarial and stress-based assessments, particularly for models intended for high-stakes deployment. These assessments should be reported alongside accuracy metrics and form the basis of model release audits.
Third, the field must shift its mindset: benchmarks are tools for assess, not endpoints. Their role is to reveal model capabilities and limitations under realistic constraints—not to be optimized at the cost of generalization.
As large multimodal models continue to advance, our ability to test them meaningfully must evolve in parallel. Stress testing offers a path forward, grounded not in artificial difficulty, but in the realities of real world.
Extended Data
@@ IV. Toward Community Standards for Evaluation Reform
Table 1: Benchmark leaderboard accuracy, (%) across representative multimodal medical benchmarks.
@@
GPT-5 67.85 58.85 87.43 86.59 80.89
Gemini-2.5 Pro 66.30 58.44 85.69 84.84 79.95
OpenAI-o3 65.63 60.08 85.49 84.75 80.89
OpenAI-o4-mini 64.08 59.47 81.43 80.46 75.91
GPT-4o 60.98 50.21 73.31 69.94 66.90
DeepSeek-VL2 45.68 40.95 54.01 38.21 33.16
LLaVA-Med-1.5 44.35 6.17 68.66
@@
GPT-5 86.59 82.91 -3.68 80.89 67.56 -13.33
Gemini-2.5 Pro 84.84 74.93 -9.91 79.95 65.01 -14.94
OpenAI-o3 84.75 82.65 -2.10 80.89 67.03 -13.86
OpenAI-o4-mini 80.46 78.44 -2.02 75.91 66.49 -9.42
GPT-4o 69.94 68.89 -1.05 66.90 37.28 -29.62
DeepSeek-VL2 38.21 32.60 -5.61 33.16 25.30 -7.86
LLaVA-Med-1.5 22.74
Notes: (i) Two-decimal precision preserved from raw logs; $\Delta$ computed as exact subtraction (two decimals). (ii) For NEJM, DeepSeek-VL2 and LLaVA-Med-1.5 lacked filtered Image+Text results (743); hence cells are “–”. If you prefer to include unfiltered NEJM entries for completeness: DeepSeek-VL2 Img+Txt=33.16 (N=947), yielding $\Delta=-7.86$ with Txt Only=25.30; LLaVA-Med-1.5 Img+Txt=29.04 (N=947), no Txt Only available.
@@
GPT-5 66.3 37.7
Gemini-2.5 Pro 67.4 37.1
OpenAI-o3 61.7 37.7
OpenAI-o4-mini 58.9 33.7
GPT-4o 46.3 3.4
@@
Table 4: Stress Test 3: Format Perturbation. Accuracy (%) on a visually-required NEJM subset before and after random reordering of answer choices. Reordering weakens shortcut patterns (e.g., fixed answer positions), reducing accuracy in the Text condition. In contrast, Image+Text performance is stable or slightly improved, suggesting vision becomes more informative when textual cues are disrupted.
@@
GPT-5 37.71 32.00 66.28 70.85
Gemini-2.5 Pro 37.14 33.14 67.42 70.28
OpenAI-o3 37.71 31.42 61.71 64.00
OpenAI-o4-mini 33.71 33.71 58.86 55.43
GPT-4o 3.43 0.57 46.29 46.85
@@
GPT-5 37.71 42.86 33.14 29.24 24.27 20.00
Gemini-2.5 Pro 37.14 45.71 38.86 35.42 26.85 17.14
OpenAI-o3 37.71 34.28 29.71 27.42 21.14 20.00
OpenAI-o4-mini 33.71 42.28 36.57 32.57 24.57 17.14
GPT-4o 3.43 4.57 1.71 2.28 0.37 2.28
@@
GPT-5 66.28 71.73 69.71 73.71 86.29 90.86
Gemini-2.5 Pro 67.43 69.71 72.00 73.71 82.29 92.57
OpenAI-o3 61.71 67.43 68.57 70.86 80.57 89.14
OpenAI-o4-mini 58.86 61.14 62.37 68.57 73.72 85.14
GPT-4o 46.29 53.71 50.28 54.28 65.14 84.00

### v3

@@ Results
Stress Test 6: Reasoning signal fidelity. We evaluated how LLMs generate and use reasoning when solving multimodal medical questions. Two complementary analyses were performed. First, we applied Chain-of-Thought (CoT) [20] prompting to 100 items from NEJM and VQA-RAD [21] to test whether explicit reasoning steps improve accuracy. Second, we manually audited model-generated explanations across benchmarks to assess their factuality, visual grounding, and consistency with the final answer. We further varied reasoning strength on OmniMedVQA [22] using OpenAI-o3 to examine whether more elaborate rationales enhance reasoning quality.
Across datasets, explicit CoT prompting yielded limited or negative gains. On NEJM and VQA-RAD, GPT-5 and Gemini-2.5 Pro changed little (0-2 pp), whereas o4-mini improved slightly (+2-5 pp). On OmniMedVQA, increasing reasoning strength produced small or inconsistent effects; longer chains sometimes increased recall but also hallucinated details.
Manual audits revealed three recurring patterns: (1) Correct answer, incorrect logic-plausible-sounding rationales containing false or fabricated visual findings. (2) Amplified visual misunderstanding-initial perceptual errors propagated through subsequent reasoning. (3) Structured but uninformative steps-syntactically coherent but clinically irrelevant statements. Examples appear in Fig. 11 and Appendix Supplementary Note 3 | More reasoning failure results.
Interpretation. Reasoning prompts do not reliably improve multimodal accuracy, and self-generated explanations often diverge from the true decision process. While larger models produce fluent, structured rationales, these frequently include non-existent image features or illogical justifications. These findings suggest that fluency in explanation does not equate to valid inference: models can produce convincing but unsupported narratives. From a translational perspective, reasoning fidelity must therefore be independently validated before such outputs can inform clinical decision support.
Table 1: Taxonomy of failure modes in health AI. Failure types are categorized by their cognitive locus, defined concisely, and linked to clinical implications and diagnostic stress tests.
Failure Category Definition Clinical Consequence Tests
Perception and Input Handling
Visual misperception Hallucinates, omits, or misinterprets image features Missed findings or false positives T2, T5
Refusal miscalibration Answers when uncertain or abstains when reliable Unsafe guesswork or care delay T2–T4
Modality neglect Disregards available image or text input Incomplete clinical synthesis T1, T2
Reasoning and Inference
Heuristic dependence Relies on superficial cues (e.g., position, phrasing) Fragile under perturbation or distribution shift T3–T5
Justification error Provides inaccurate or unsupported rationale Misleading interpretability T2, T6
Logical inconsistency Internal conflict across input, output, or reasoning Reduced trust and auditability T1–T6
Output and Communication
Fluent factual error Persuasive language masks incorrect claims Diagnostic or therapeutic misguidance T2, T6
Unsafe recommendation Suggests harmful or incomplete action plans Clinical risk or treatment harm T5, T6
Taxonomy of failure modes
Across the six stress tests, recurring limitations cluster into three stages of model behavior:input handling, reasoning and inference, and output communication. Each represents a distinct reliability dimension that links model behavior to potential sources of clinical risk. A concise summary is shown below; the full taxonomy, including cognitive mechanisms and harm potential, appears in - Data Table 1.
This taxonomy organizes observed failure patterns along the cognitive pathway from perception to communication, providing a structured lens for assessing readiness, safety, and interpretability in health AI models.
Benchmarking the Benchmarks: What Are We Really Measuring?
@@
Despite similar leaderboard scores, Section Stress testing exposes robustness gaps in LLMs revealed substantial variation in how models respond to medical stressors-ranging from hallucinated justifications to modality overreliance and format-based shortcuts. These divergent behaviors raise a fundamental question:
@@
If models succeed on benchmarks but fail under stress, what do these benchmarks actually measure-and how can we improve them?
@@
Motivation: Stress Test Failures Reveal Benchmark Gaps
The stress tests in Section Stress testing exposes robustness gaps in LLMs revealed that high benchmark scores do not guarantee robust model behavior. Models that performed well under standard conditions often failed when subjected to small perturbations-such as removing the image, reordering answer options, replacing distractors, or introducing misleading visual inputs.
Critically, these failure modes varied by benchmark. On NEJM, removing visual input led to steep performance drops, suggesting strong dependence on image understanding. In contrast, scores on JAMA remained stable-indicating that many items could be answered from text alone. Similarly, while chain-of-thought prompting provides explicit reasoning scaffolding that benefits distilled models such as o4-mini, its impact on NEJM was substantially greater than on VQA-RAD, underscoring the heterogeneous reasoning demands across benchmarks. In Test 5, even subtle visual substitutions caused top models to confidently choose incorrect answers-highlighting shortcut reliance on visual-answer associations in some datasets but not others.
@@
Clinician-Informed Benchmark Profiling
To better understand what each benchmark actually tests, we developed a structured, clinician-guided rubric spanning ten medically meaningful dimensions. These axes capture key demands surfaced by our stress tests-ranging from reasoning complexity and medical context to uncertainty handling, visual detail, and multi-view alignment. The full rubric is shown in Fig 4.
Each benchmark was independently annotated by three board-certified clinicians per axis, using a 3-point ordinal scale. We report the median score for each benchmark-axis pair, and measure inter-annotator agreement using Fleiss’ $\kappa$ [28]. Agreement was moderate to strong across axes (range: 0.67-0.90). Highest agreement was observed on binary-seeming attributes such as Text-only Solvable ($\kappa$ = 0.90) and Clinical Context ($\kappa$ = 0.86); lower agreement appeared on more subjective dimensions such as Visual Detail Required ($\kappa$ = 0.67).
Figure 4: Benchmarks differ in reasoning and visual complexity. a, Clinician-labeled profiles reveal that widely used multimodal health AI benchmarks vary substantially in the reasoning and visual demands they place on models. Each dataset is shown using a ripple marker, where the central point denotes the mean complexity rating and the thin concentric ring encodes variability across annotators. b, Benchmarks ranked by average reasoning complexity (with variability). c, Benchmarks ranked by average visual complexity (with variability). These structural differences help explain the benchmark-specific fragilities observed in Section Stress testing exposes robustness gaps in LLMs, and highlight the risk of treating all benchmarks as interchangeable measures of readiness.
Visualizing the Benchmark Landscape
@@
  • VQA-RAD, PMC-VQA and MIMIC-CXR [23] are visually dependent but low in inference complexity.
@@
This landscape helps to explain the benchmark-specific failures observed in Section Stress testing exposes robustness gaps in LLMs. For instance, models trained on VQA-RAD may excel at image localization but struggle on diagnosis-oriented tasks like NEJM, which require cross-modal reasoning and medical inference.
Implications for Model Evaluation and Design
These findings underscore the need to treat benchmarks not as monolithic performance targets, but as diagnostic tools that reveal distinct capabilities. We outline four implications:
@@
  • Evaluation protocols should routinely incorporate adversarial and stress-based assessments. These assessments should be reported alongside accuracy metrics and form the basis of model release audits, particularly for models intended for high-stakes deployment.
@@
Discussion
@@
We have shown that current benchmarks often overstate model readiness. Despite impressive leaderboard scores, many leading models exhibit inconsistent behavior, reliance on superficial patterns, and fragile reasoning under mild input perturbations. These findings challenge the prevailing assumption that benchmark success signals medical viability.
@@
Today’s benchmarks offer a convenient but incomplete lens into model capability. Their design often emphasizes answer correctness while ignoring whether the answer is reached through medically valid reasoning, multimodal understanding, or robust generalization. As demonstrated in our stress tests, models can perform well by leveraging shallow cues, such as positional biases in answer options or memorized response formats, without engaging meaningfully with the underlying medical content. Such behaviors are not easily detected by conventional metrics. For example, reasoning models may improve benchmark scores while still producing hallucinated or implausible rationales. Reinforcement learning approaches may optimize for token-level reward signals rather than faithful reasoning. The result is a misleading sense of progress: performance improves numerically, but not behaviorally.
@@
Medical readiness is a multidimensional construct. In real-world settings, models must tolerate missing or noisy data, justify their decisions in a manner clinicians can understand, and reason across time, modality, and context. Performance must be not only accurate but also reliable, interpretable, and safe under uncertainty. Unlike domains such as mathematics or programming, where ground truth is unambiguous and logic is formal, healthcare requires flexible, contextual reasoning. Attempts to port reasoning strategies from other domains, such as multi-agent planning or chain-of-thought prompting, often fall short in medicine, where ambiguity, incomplete information, and medical nuance are the norm. Without alignment to this reality, advances in model capability risk being confined to artificial test settings.
@@
Static benchmark scores provide little insight into how models behave under real-world uncertainty. To address this limitation, we propose structured stress testing as a core component of evaluation. As shown in Section 2, our tests T1–T5 revealed distinct vulnerabilities, models often succeed for the wrong reasons, retaining high accuracy even when visual input is removed (T1–T2, see Example 9-11); they show brittle behavior under minor perturbations such as distractor reordering, weakening, or visual substitution (T3–T5, See Example 9-14); and they frequently produce fabricated reasoning chains, generating confident but incorrect explanations that mask underlying shortcut behaviors (T6, See Example 15-17). These fragilities were hidden in leaderboard performance alone. We generalize this approach into a framework of modular tests, each targeting a specific vulnerability such as neglect of visual input, dependence on spurious answer patterns, or overconfident reasoning when evidence is absent. By isolating these behaviors, stress tests enable systematic comparison across models and tasks, offering a principled means of diagnosing brittleness that static benchmarks obscure. In doing so, they establish a foundation for evaluation protocols that move beyond leaderboards and toward measures of robustness and trustworthiness.
We propose several actionable recommendations for the field: First, benchmarks should be accompanied by metadata that characterizes their reasoning and visual complexity, using structured rubrics like those presented in Section Benchmarking the Benchmarks: What Are We Really Measuring?. This allows model performance to be interpreted in medically meaningful dimensions rather than aggregated scores alone. Second, evaluation protocols should routinely incorporate adversarial and stress-based assessments, particularly for models intended for high-stakes deployment. These assessments should be reported alongside accuracy metrics and form the basis of model release audits. Third, the field must shift its mindset: benchmarks are tools for assess, not endpoints. Their role is to reveal model capabilities and limitations under realistic constraints, not to be optimized at the cost of generalization. As large multimodal models continue to advance, our ability to test them meaningfully must evolve in parallel. Stress testing offers a path forward, grounded not in artificial difficulty, but in the realities of real world.
Supplementary Information for:
The Illusion of Readiness in Health AI
Supplementary Figures
Figure 5: Illustration of model behavior under Stress Tests 2 and 3: Case 1. We apply three modifications to the original image-based multiple-choice question (ground truth: Dermatomyositis): (i) remove the image, (ii) shuffle the answer options, and (iii) replace one option with “Unknown.” When the image is absent, the appropriate behavior is to abstain (or to select “Unknown” if available). However, most models still provide a diagnosis, often repeating Dermatomyositis even after the options are shuffled, whereas GPT-4o is the only model that consistently refuses to answer without the image.
Figure 6: Illustration of model behavior under Stress Tests 2 and 3: Case 2. We apply three modifications to the original image-based multiple-choice question (ground truth: Chiari I malformation): (i) remove the image, (ii) shuffle the answer options, and (iii) replace one option with “Unknown.” When the image is absent, the expected behavior is to abstain (or to select “Unknown” if available). Nevertheless, most models continue to provide a diagnosis, often repeating the original ground truth, whereas GPT-4o is the only model that consistently either abstains or selects “Unknown” in the absence of the image.
Figure 7: Illustration of model behavior under Stress Tests 2 and 3: Case 3. We apply three modifications to the original image-based multiple-choice question (ground truth: Pneumatosis cystoides coli): (i) remove the image, (ii) shuffle the answer options, and (iii) replace one option with “Unknown.” When the image is absent, the expected behavior is to abstain (or to select “Unknown” if available). Nevertheless, most models continue to provide a diagnosis, often repeating the original ground truth, whereas GPT-4o is the only model that consistently refuses to answer without the image.
Supplementary Tables
@@
Table 2: Benchmark leaderboard accuracy (%) across representative multimodal medical benchmarks.
@@
GPT-5 67.85 [63.41,72.06] 58.85 [54.53,63.17] 87.43 [84.53,90.14] 86.59 [84.66,88.52] 80.89 [78.06,83.71]
Gemini-2.5 Pro 66.30 [61.86,70.51] 58.44 [54.12,62.76] 85.69 [82.59,88.59] 84.84 [82.82,86.85] 79.95 [77.12,82.77]
OpenAI-o3 65.63 [61.20,70.07] 60.08 [55.76,64.40] 84.91 [81.82,88.01] 84.75 [82.65,86.85] 80.89 [78.06,83.71]
OpenAI-o4-mini 64.08 [59.65,68.51] 59.47 [55.14,63.79] 81.43 [78.14,84.72] 80.46 [78.09,82.73] 77.66 [74.56,80.62]
GPT-4o 60.98 [56.54,65.41] 52.26 [47.74,56.79] 74.08 [70.21,77.76] 69.94 [67.22,72.57] 64.33 [60.83,67.83]
DeepSeek-VL2 45.68 [41.02,50.11] 38.89 [34.57,43.21] 54.16 [49.90,58.41] 38.21 [35.85,41.37] 34.59 [31.22,38.09]
LLaVA-Med-1.5 46.34 [41.69,51.00] 6.17 [4.12,8.44] 68.67 [64.80,72.73] 22.83 [20.44,25.31] 29.34 [26.11,32.71]
@@
GPT-5 86.59 [84.66,88.52] 82.91 [80.72,85.10] -3.68 80.89 [78.06,83.71] 67.56 [64.20,70.93] -13.33
Gemini-2.5 Pro 84.84 [82.82,86.85] 74.93 [72.39,77.48] -9.91 79.95 [77.12,82.77] 65.01 [61.51,68.37] -14.94
OpenAI-o3 84.75 [82.65,86.85] 82.65 [80.46,84.84] -2.10 80.89 [78.06,83.71] 67.03 [63.53,70.39] -13.86
OpenAI-o4-mini 80.46 [78.09,82.73] 78.44 [76.07,80.81] -2.02 77.66 [74.56,80.62] 66.49 [63.12,69.85] -11.17
GPT-4o 69.94 [67.22,72.57] 68.89 [66.17,71.52] -1.05 64.33 [60.83,67.83] 37.28 [33.78,40.78] -27.05
DeepSeek-VL2 38.21 [35.85,41.37] 32.60 [29.97,35.32] -5.61 34.59 [31.22,38.09] 25.30 [22.34,28.53] -9.29
LLaVA-Med-1.5 22.83 [20.44,25.31] 29.34 [26.11,32.71]
Notes: (i) Two-decimal precision preserved from raw logs; $\Delta$ computed as exact subtraction (two decimals). (ii) LLaVA-Med-1.5 lacked Txt Only results, and hence cells are “–”.
@@
GPT-5 66.29 [59.43,73.14] 37.71 [30.29,45.14]
Gemini-2.5 Pro 67.43 [60.00,74.29] 37.14 [30.29,44.57]
OpenAI-o3 61.71 [54.86,68.57] 37.71 [30.29,45.14]
OpenAI-o4-mini 58.86 [51.43,66.29] 33.71 [26.86,40.57]
GPT-4o 46.29 [38.86,53.71] 3.43 [1.14,6.29]
@@
Table 5: Stress Test 3: Format Perturbation. Accuracy (%) on a visual-required NEJM subset before and after random reordering of answer choices. Reordering weakens shortcut patterns (e.g., fixed answer positions), reducing accuracy in the Text condition. In contrast, Image+Text performance is stable or slightly improved, suggesting vision becomes more informative when textual cues are disrupted.
@@
GPT-5 37.71 [30.29,45.14] 32.00 [25.14,38.86] 66.29 [59.43,73.14] 70.86 [64.00,77.71]
Gemini-2.5 Pro 37.14 [30.29,44.57] 33.14 [26.29,40.00] 67.43 [60.00,74.29] 70.29 [62.86,77.14]
OpenAI-o3 37.71 [30.29,45.14] 31.43 [24.57,38.29] 61.71 [54.86,68.57] 64.00 [56.57,70.86]
OpenAI-o4-mini 33.71 [26.86,40.57] 33.71 [26.86,40.57] 58.86 [51.43,66.29] 55.43 [48.00,62.86]
GPT-4o 3.43 [1.14,6.29] 0.57 [0.00,1.71] 46.29 [38.86,53.71] 46.86 [39.43,54.29]
@@
GPT-5 37.71 [30.29,45.14] 42.86 [35.43,50.29] 33.14 [26.29,40.00] 29.14 [22.29,36.00] 24.57 [18.29,30.86] 20.00 [14.29,26.29]
Gemini-2.5 Pro 37.14 [30.29,44.57] 45.71 [38.29,53.14] 38.86 [32.00,46.29] 35.43 [28.57,42.29] 26.86 [20.57,33.71] 17.14 [12.00,22.86]
OpenAI-o3 37.71 [30.29,45.14] 34.29 [27.43,41.71] 29.71 [22.86,36.57] 27.43 [21.14,34.29] 21.14 [15.43,27.43] 20.00 [14.29,26.29]
OpenAI-o4-mini 33.71 [26.86,40.57] 42.29 [34.86,49.14] 36.57 [29.71,43.43] 32.57 [25.71,39.43] 24.57 [18.29,31.43] 17.14 [12.00,22.86]
GPT-4o 3.43 [1.14,6.29] 4.57 [1.71,8.00] 1.71 [0.00,4.00] 2.29 [0.57,4.57] 0.57 [0.00,1.71] 2.29 [0.57,4.57]
@@
GPT-5 66.29 [59.43,73.14] 71.43 [64.57,77.71] 69.71 [62.86,76.00] 73.71 [66.86,80.00] 86.29 [81.14,90.86] 90.86 [86.29,94.86]
Gemini-2.5 Pro 67.43 [60.00,74.29] 69.71 [62.86,76.69] 72.00 [65.14,78.29] 73.71 [66.86,80.00] 82.29 [76.57,87.43] 92.57 [88.57,96.00]
OpenAI-o3 61.71 [54.86,68.57] 67.43 [60.57,74.29] 68.57 [61.71,75.43] 70.86 [64.00,77.14] 80.57 [74.29,86.29] 89.14 [84.57,93.71]
OpenAI-o4-mini 58.86 [51.43,66.29] 61.14 [53.71,68.61] 61.71 [54.84,68.57] 68.57 [61.71,75.43] 73.71 [66.86,80.00] 85.14 [80.00,90.29]
GPT-4o 46.29 [38.86,53.71] 53.71 [46.29,61.14] 50.29 [42.86,57.71] 54.29 [46.86,61.71] 65.14 [57.71,72.00] 84.00 [78.29,89.14]